OpenStudy (anonymous):

If the diameter of the base remains constant, by what factor should the altitude be multiplied to produce a cone with twice volume as the original?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

1/2?

OpenStudy (anonymous):

how did you get it?

OpenStudy (amistre64):

how do we determine the volume of a cone?

OpenStudy (anonymous):

V= 1/3 pi r^2 h

OpenStudy (amistre64):

now multiply both sides by 2

OpenStudy (anonymous):

2V =2(1/3 pi r^2h)

OpenStudy (amistre64):

notice that r doesnt change, 1/3 and pi dont change ... we only change the height what did we alter?

OpenStudy (amistre64):

2V = 1/3 pi r^2 2h

OpenStudy (anonymous):

you said that i will multiple the other side means: 2/3*2pi*2r^2*2h = 2V

OpenStudy (amistre64):

no .... you cant make up your own math rules .... 2(ab) does not equate to 2a2b

OpenStudy (amistre64):

if a remains the same, then we must have altered b .. by 2

OpenStudy (anonymous):

what do you mean, 1/3 *pi*r^2 are a fixed value?

OpenStudy (amistre64):

the formula for volume doesnt change does it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

can you show me your solutions

OpenStudy (amistre64):

if the formula is consistent for all cones ... then V = 1/3 pi r^2 h is for any size cone since 1/3 and pi are constant for any size cone ... and we are told ... i did show you, they are posted above

OpenStudy (amistre64):

why do i bother posting this stuff if noone has the gumption to just scroll up to look at it ....

OpenStudy (anonymous):

i am looking

OpenStudy (amistre64):

:) thnx

OpenStudy (anonymous):

no one wants to answer my qs so im looking at others while i wait and this is intresting

OpenStudy (anonymous):

so that means that if we double the height, it doubles the volume?

OpenStudy (amistre64):

V1 the original cone: 1/3 pi r^2 h1 V2 the scaled cone: 1/3 pi r^2 h2 V2 = 2V1 1/3 pi r^2 h2 = 2(1/3 pi r^2 h1) factor out common factors h2 = 2 h1

OpenStudy (anonymous):

so the golden rule what you do to one side do to the other. and we are inding x witch is ether our voluem or the fraction we want

OpenStudy (amistre64):

we double the height to double the volume which is logically what we would do if this is a cylindar ... and a cone is just 1/3 the colume of its representative cylindar

OpenStudy (anonymous):

that make sense ...lol...you must do it earlier so that we can move on another problem

OpenStudy (anonymous):

ok the h thing makes sense because 2 x 1 right?(sorry if im bothering you